OpenStudy (anonymous):

@aajugdar Solve (z + 6)^2 = 5.

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

hmm so you can write (z+6)^2=5 as \[z^2+12z+36=5 \]

OpenStudy (anonymous):

now \[z^2+12z+31=0 \]

OpenStudy (anonymous):

this one cant be factorised so we use the formula \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2 }\]

OpenStudy (anonymous):

2a* in denominator

OpenStudy (anonymous):

Im 100% confused

OpenStudy (anonymous):

^ there you go

OpenStudy (anonymous):

This would be the answer.

OpenStudy (anonymous):

no there is minus sign b4 6

OpenStudy (anonymous):

\[x=-6\pm \sqrt{5}\]

OpenStudy (anonymous):

Yeah sorry I sent the wrong thing. Thank you. You're really smart.

OpenStudy (anonymous):

haha,i dnt deserve that compliment :P

OpenStudy (anonymous):

You're the reason I passed the test. So yah you do LOL

OpenStudy (anonymous):

LOL